Two samples of n = 5 subjects were analyzed with an independent-samples t-test, producing a t = 2.295.

a) What is the tcrit ?

b) Does this indicate a significant difference in a two-tailed test with α = .05?
Answer #1

The null and alternative hypothesis is:

H0: µ12

Ha: µ1 µ2

The degrees of freedom = n1+n2 -2 =5 + 5-2= 8

a) The critical value at 5% significance level and 8 degrees of freedom is, tcrit= 2.306

b) The value of t test statistic is less than critical value of t. Hence we accept null hypothesis at 5% significance level and conclude that there is no significant difference.